OpenStudy (anonymous):

What is the equation of a line perpendicular to the line y = 4/9x - 2 that passes through the point (4,3)? Options: y = 9/4x + 6 y = -4/9x - 2 y = -9/4x+ 12 y = -9/4 - 12

OpenStudy (zale101):

the slope equation is y = 4/9x - 2 as in gradient intercept form. (y=mx+b) after u have your slope formula, do the point-slope formula y-y1=m(x-x1)

OpenStudy (anonymous):

... I don't really get how to turn it into the point slope formula?

OpenStudy (zale101):

(4,3) is (x1,y1) 4/9 is the slope

OpenStudy (zale101):

plug in into this formula: y-y1=m(x-x1)

OpenStudy (anonymous):

I just don't get this really and how it relates to the problem...

OpenStudy (zale101):

the equation of a straight line in the form y − y1 = m(x − x1) where m is the slope of the line and (x1, y1) are the coordinates of a given point on the line

OpenStudy (anonymous):

I'm sorry your just not really helping me...

OpenStudy (anonymous):

I'm not following what you're trying to say..

OpenStudy (zale101):

what part u did't get, ur confusing me?

OpenStudy (anonymous):

I don't get how all this is helping me find the line perpendicular to my original line.

OpenStudy (zale101):

A line perpendicular to y = 4/9x - 2 it's written as y=mx+b where m=slope, b=y-intercept so has a slope (m) of 4/9 using point (4, 3): (y - y1) = m(x - x1) (y - 3) = (4/9)(x - 4)

OpenStudy (zale101):

now simplify

OpenStudy (anonymous):

I'm sorry just the way you're saying it isn't making sense at all. I think I just need someone else helping on this one.

OpenStudy (zale101):

ok, good luck